OpenStudy (eric_d):

g(x)=(x-2)(x-4), x bigger or equal than 3 Find inverse function

OpenStudy (ikram002p):

g(x)=(x-2)(x-4) y=x^2-6x+8 now complete a a square

OpenStudy (ikram002p):

y +\((\frac{-6}{2})^2\) =x^2-6x+8 +\((\frac{-6}{2})^2\)

OpenStudy (eric_d):

(x+9)^2-1

OpenStudy (ikram002p):

mmm u made a small mistake

OpenStudy (ikram002p):

when u add (6/2)^2 y=x^2-6x+8+(6/2)^2-(6/2)^2 y=x^2-6x+ (3)^2-(3)^2 +8 y=x^2-6x+ (3)^2-9 +8 so it will be :- y=x^2-6x+ (3)^2-1 y=(x-3)^2 -1

OpenStudy (ikram002p):

@eric_d got it ?

OpenStudy (eric_d):

okay, jst realised

OpenStudy (ikram002p):

ok so nw can u find the inverse ?

OpenStudy (eric_d):

I don't know

OpenStudy (ikram002p):

ohh its ok :D so u have y=(x-3)^2-1 the inverse means to write the function with respect to y that means solve for x ( try to make x in one side by its own and the other side includs y)

OpenStudy (eric_d):

ok

OpenStudy (eric_d):

understood

OpenStudy (ikram002p):

y=(x-3)^2-1 step 1:- y+1=(x-3)^2 stwp 2 :- \(\pm \sqrt{y+1}= x-3 \)

OpenStudy (eric_d):



OpenStudy (eric_d):

That will be the final answer rite

OpenStudy (eric_d):

@ikram002p

OpenStudy (ikram002p):

yeah right !

OpenStudy (eric_d):

Thanks @ikram002p

OpenStudy (ikram002p):

np :)

OpenStudy (eric_d):

Can u fan me So, I can message u.... @ikram002p

OpenStudy (ikram002p):

done :D

OpenStudy (eric_d):

Thanks